Graph the function y=x^2 -6x+1
A. State the domain over which the function is increasing.
B. state the range of the function.
C. Describe the end behavior of the function.

Answer by ewatrrr(24785)


Hi
y= x^2 -6x+1 |Parabola opening upward 1 >0, y-intercept P(0,1)
y = (x-3)^2 - 9 + 1
y = (x-3)^2 - 8 V(3,-8), x = 3 line of symmetry
A. State the domain over which the function is increasing. x > 3
B. state the range of the function. y ≥ -8
C. Describe the end behavior of the function. Infinity